OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of 9 cm and 12 cm

OpenStudy (anonymous):

Use Pythagorean Theorem; \[a ^{2}+b ^{2}=c ^{2}\] So.. \[9^{2}+12^{2}=c ^{2}\] \[9^{2}=9*9=81\] \[12^{2}=12*12=144\] Which means; \[81+144=c ^{2}\] \[225=c^{2}\] \[\sqrt{225}=c\] \[15=c\] So, 15 is the hypotenuse.

OpenStudy (anonymous):

81 + 144 =225. SQRT of 225 = 15 C=15

OpenStudy (anonymous):

o.o that looks confusing but thank you SO much your amazing :]

OpenStudy (anonymous):

It usually looks harder than it is, haha. If you need anymore help I'll be glad to try and do so. :D

OpenStudy (anonymous):

thank you:] hahah im horrible in math